# Better Fair than Sorry: Adversarial Missing Sensitive Value Imputation for Fair GNNs

### v2

@@ Abstract
This paper addresses the problem of learning fair Graph Neural Networks (GNNs) under missing protected attributes. GNNs have achieved state-of-the-art results in many relevant tasks where decisions might disproportionately impact specific communities. However, existing work on fair GNNs assumes that either protected attributes are fully observed or that the missing protected attribute imputation is fair. In practice, biases in the imputation will be propagated to the model outcomes, leading them to overestimate the fairness of their predictions. We address this challenge by proposing Better Fair than Sorry (BFtS), a fair missing data imputation model for protected attributes used by fair GNNs. The key design principle behind BFtS is that imputations should approximate the worst-case scenario for fairness—i.e. when optimizing fairness is the hardest. We implement this idea using a 3-player adversarial scheme where two adversaries collaborate against a GNN-based classifier to maximize its bias. Experiments using synthetic and real datasets show that BFtS often achieves a better fairness $\times$ accuracy trade-off than existing alternatives.
@@ 1 Introduction
With the increasing popularity of machine learning models for high-stakes decision-making, it has become a consensus that (1) these models carry implicit biases and (2) these biases should be addressed to improve the fairness of algorithmic decisions Ghallab (2019). The disparate treatment of such models towards African Americans and women has been illustrated in the well-documented COMPAS Angwin et al. (2022) and Apple credit card Vigdor (2019) cases. While there has been extensive research on fair algorithms and fair machine learning, the proposed solutions have mostly disregarded important challenges that arise in real-world settings.
@@
Figure 1: Motivation: In Figure (a), a graph machine learning algorithm is applied to decide who receives credit (positive or negative) based on a possibly missing sensitive attribute, gender (and binary only for illustrative purposes). As shown in Figure (b), traditional missing data imputation does not account for outcomes (positive/negative) and thus their imputed values can under-represent the bias of the complete dataset—demographic parity (DP) is $0.09$ in this example. This paper proposes BFtS, an imputation method for graph data that optimizes fairness in the worst-case imputation scenario using adversarial learning, as shown in Figure (c), where DP is $0.47$.
@@
Existing work in fair machine learning has focused on tabular, image, and text data Mehrabi et al. (2021); Pessach and Shmueli (2022). However, in several applications, data can be naturally modeled as graphs (or networks), representing different objects, their relationships, and attributes Barabási (2016); Newman (2018). Graph Neural Networks (GNNs) have achieved state-of-the-art results in many graph machine learning tasks, including node classification, link prediction, and graph classification Gilmer et al. (2017); Kipf and Welling (2017); Veličković et al. (2017a); Hamilton et al. (2017). For instance, in the case of link prediction in a professional network, such as LinkedIn, which is a key recruiting and networking tool, link recommendations should not be biased against protected groups Fabbri et al. (2020); Karimi et al. (2018); Current et al. (2022). However, guaranteeing fairness in graph data is a challenge due to well-known correlations in the network caused by homophily and influence. Anagnostopoulos et al. (2008); Aral et al. (2009).
@@
An additional challenge for fair machine learning is the pervasiveness of missing data in many applications Little and Rubin (2019); Buck (1960); Enders (2010). This is particularly critical because fair algorithms often require knowledge of protected (or sensitive) attributes that are more likely to be missing due to biases in the collection process or privacy concerns (e.g., gender, race, age). As an example, data absence in census surveys is correlated with sensitive attributes such as race, etc. O’Hare and O’Hare (2019); Warren and Passel (1987). Similarly, gender information is often missing in health surveys Weber et al. (2021); Molenberghs and Kenward (2007). Missing data is also a concern in networked data, especially when tracking the transmission of diseases Huisman (2009); Gile and Handcock (2006, 2017); Ghani et al. (1998). As a consequence, many fair machine learning algorithms require the use of missing imputation methods to mitigate their biases, and biases in imputed values might be inadvertently propagated to the final decisions Mansoor et al. (2022).
@@
This paper addresses the problem of mitigating bias in graph machine-learning algorithms in the presence of missing sensitive attributes. We propose an adversarial imputation method that generates challenging instances for a broad class of fair algorithms based on GNNs. Figure 1 motivates our approach. We use a machine learning algorithm that decides whether individuals should or should not receive credit. Figure 0(a) shows both the gender and outcome of each individual. The genders of some samples are unknown in Figure 0(a) and the demographic parity (DP) considering only the observed values is $0.25$ (DP is a bias metric described in Section 3.3). We illustrate how different missing data imputations affect the fairness of credit decisions in Figures 0(b) and 0(c). A straightforward imputation is demonstrated in Figure 0(b), where the gender of the majority of neighbors is assigned to the missing attribute. This results in a best-case scenario in terms of fairness with a DP of $0.09$. However, in the worst-case scenario, shown in Figure 0(c), DP is $0.47$. The bias of the dataset might be highly underrepresented in the imputation of Figure 0(b). The final decision of a fair algorithm will still be unfair if the imputation underestimates the bias. Our empirical results provide strong evidence that off-the-shelf imputation algorithms often misguide fair graph machine learning algorithms due to underestimated values of bias.
@@
We propose Better Fair than Sorry (BFtS), which is a 3-player adversarial learning framework for missing data imputation that produces worst-case imputations for a fair GNN. Figure 2 provides an overview of the proposed approach. The GNN $f_{C}$ interacts with two adversaries, $f_{A}$ and $f_{I}$. $f_{A}$ attempts to predict sensitive values based on the representations learned by $f_{C}$ while $f_{I}$ imputes missing values that help $f_{A}$ maximize its accuracy. As high accuracy of $f_{A}$ indicates high bias in the representations learned by $f_{C}$, $f_{C}$’s parameters are optimized to minimize the bias of its representations in the worst-case in terms of fairness. Our experiments show that BFtS achieves a better accuracy $\times$ fairness trade-off than existing approaches using real datasets and synthetic datasets.
We summarize the contributions of this paper as follows:
  • We investigate a novel problem: fair imputation of missing sensitive data for graph machine learning.
@@
  • We propose Better Fair than Sorry (BFtS), a novel 3-player adversarial learning framework for the imputation of missing sensitive data that produces challenging values for graph-based fairness.
  • We conduct a theoretical analysis showing that BFtS is more robust to convergence issues in training.
@@
  • We demonstrate that BFtS achieves a better fairness $\times$ accuracy trade-off than alternative solutions.
@@ 1.1 Related Work
In this section, we will overview related work regarding missing data imputation and fairness in graph machine learning.
1.1.1 Fairness in graph machine learning
@@
We divide the work in this area into three main directions (pre-processing, extended objective functions, and adversarial learning). FairOT and FairDrop are pre-processing techniques that remove the bias in the graph or the node attributes before training the GNN Laclau et al. (2021); Spinelli et al. (2021). Fairwalk, Crosswalk, Debayes MONET, NIFTY, FairVGNN, and FairGAE modify the objective function of GNNs to learn fair representations during training Rahman et al. (2019); Khajehnejad et al. (2022); Buyl and De Bie (2020); Palowitch and Perozzi (2019); Agarwal et al. (2021); Wang et al. (2022); Fan et al. (2021). CFC, FLIP, DKGE, and Debias train adversarial networks along with GNN for fair prediction Bose and Hamilton (2019); Masrour et al. (2020); Arduini et al. (2020); Zhang et al. (2018). However, notice that different from our work, all these methods require fully observed sensitive attributes to ensure fairness.
1.1.2 Missing Data Imputation and fairness
@@
Missing data can be imputed using unconditional mean, reconstruction, preferential attachment, autoencoders, etc. Donders et al. (2006); Pereira et al. (2020); Huisman (2009). The traditional procedure for handling missing data is to first impute the missing values, and then solve the task Rossi et al. (2021); Buck (1960). SAT is a matching-based GNN for graphs with missing attributes Chen et al. (2020) but it does not account for fairness. Training a classifier from imputed data can amplify the unfairness of a machine learning model, as discussed in Subramonian et al. (2022); Zhang and Long (2021); Feng et al. (2023). Some studies try to generate fair feature imputations. Subramonian et al. (2022) has modified mean aggregation feature imputation to minimize discrimination risk against protected groups. Feng et al. (2023) has proposed recursive partitioning and fair uniform resampling for fair classification with missing values. However, as the approaches discussed earlier, these methods require full knowledge of the protected attributes.
1.1.3 Fairness with Missing Protected attributes
@@
These studies are the ones most related to our work. One extreme assumption is the complete unavailability of sensitive information. For instance, Hashimoto et al. (2018) takes into account the worst-case distribution across a given group size, and Lahoti et al. (2020) adversarially reweighs training samples to optimize the average loss. Chai and Wang (2022) employs top $k$ average loss minimization to guarantee fairness. However, in practice, some sensitive information is often available and relatively little information can also significantly improve fairness Chai and Wang (2022). A more practical approach is, therefore, assuming that limited sensitive information is available. FairGNN Dai and Wang (2021), which is the most related work to our work, learns fair Graph Neural Network (GNN) representations while handling missing sensitive attributes in the context of node classification. The key differences between our approach (BFtS) and FairGNN are the following: (1) we propose a fair missing protected value imputation method using a 3-player scheme to generate worst-fairness imputations; (2) we propose using the LDAM loss Cao et al. (2019) instead of the cross-entropy loss to handle imbalances in the data; and (3) we consider a more comprehensive set of datasets in our evaluation. Nguyen et al. (2017); Vandenhende et al. (2019) have applied 3-player adversarial networks to improve classification and the stability of adversarial learning. We also compare our method against RNF Du et al. (2021), which is a more recent approach that generates proxy annotations for missing sensitive attributes by training the classification network with generalized cross-entropy loss Zhang and Sabuncu (2018). Our experiments, based on four real and one simulation datasets, show that BFtS often outperforms FairGNN and RNF in terms of the fairness $\times$ accuracy trade-off.
@@ 2 Fairness-Aware Adversarial Missing Data Imputation for Graphs
In this section, we introduce a graph learning framework, which we call BFtS (Better Fair than Sorry), that allows fair GNN training in a settings with missing sensitive data. BFtS applies a 3-player adversarial learning scheme, where the fair GNN is trained jointly with two adversaries—one to predict sensitive attributes based on GNN embeddings and another to impute missing values that minimize fairness. As a result, fairness assessments by the GNN are based on the worst-case imputed values generated by the adversarial imputation.
2.1 Notation
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X})$ be an undirected graph where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, and $\mathcal{X}$ are node attributes. The matrix $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix of $\mathcal{G}$ where $A_{uv}=1$ if there is an edge between nodes $u$ and $v$ and $A_{uv}=0$, otherwise. Let the sensitive (or sensitive) attribute be $s$. We assume the semi-supervised scenario, where $\mathcal{V}_{L}\subseteq\mathcal{V}$ nodes are labeled, with binary labels $y_{v}\in\{0,1\}$. We focus on the setting where sensitive attributes might be missing and only $\mathcal{V}_{S}\subseteq\mathcal{V}_{L}$ nodes include the information of the sensitive attribute $s_{v}$ for each node $v$. To simplify the discussion, we will assume that $s_{v}\in\{0,1\}$. Finally, we will consider the node classification task, where each node $v$ has a class label $y_{v}$. We notice that all these assumptions have as their main purpose to facilitate exposition and can be generalized.
2.2 Graph Neural Network
Two of the models (or players) learned by BFtS are GNNs. Without loss of generality, we will assume that they are message-passing neural networks Gilmer et al. (2017), which we will simply refer to as GNNs. A GNN learns node representations $\mathbf{h}_{v}$ for each node $v$ that can be used to predict (categorical or continuous) values $y_{v}=f(\mathbf{h}_{v})$. These representations are learned via successive $\mathsf{AGGREGATE}$ and $\mathsf{COMBINE}$ operations defined as follows:
$\mathbf{h}_{v}^{k}\!=\!\mathsf{COMBINE}^{k}(\mathbf{h}_{v}^{k-1},\mathsf{AGGREGATE}^{k-1}({\mathbf{h}_{u}^{k-1},A_{u,v}\!=\!1}))$
Figure 2: Proposed 3-player framework for fair GNN training with missing data imputation (BFtS). $f_{C}$ generates node representations by minimizing the classification loss $\mathcal{L}_{C}$ (Eqn. 1) and maximizing sensitive attribute prediction loss $\mathcal{L}_{A}$ (Eqn. 3). $f_{A}$ predicts sensitive attributes using the representations from $f_{C}$ by minimizing $\mathcal{L}_{A}$. $f_{I}$ predicts missing values by minimizing the imputation loss $\mathcal{L}_{I}$ (Eqn. 2) and maximizing $\mathcal{L}_{A}$. $\hat{y}$, $\hat{si}$, $\hat{sa}$ are the predictions of $f_{C}$, $f_{I}$ and $f_{A}$, respectively.

### v3

@@ Abstract
Graph Neural Networks (GNNs) have achieved state-of-the-art results in many relevant tasks where decisions might disproportionately impact specific communities. However, existing work on fair GNNs often assumes that either protected attributes are fully observed or that the missing protected attribute imputation is fair. In practice, biases in the imputation will propagate to the model outcomes, leading them to overestimate the fairness of their predictions. We address this challenge by proposing Better Fair than Sorry (BFtS), a fair missing data imputation model for protected attributes. The key design principle behind BFtS is that imputations should approximate the worst-case scenario for fairness—i.e. when optimizing fairness is the hardest. We implement this idea using a 3-player adversarial scheme where two adversaries collaborate against a GNN-based classifier, and the classifier minimizes the maximum bias. Experiments using synthetic and real datasets show that BFtS often achieves a better fairness $\times$ accuracy trade-off than existing alternatives.
@@ 1 Introduction
With the increasing popularity of machine learning models for high-stakes decision-making, it has become a consensus that these models carry implicit biases, and biases should be addressed to improve the fairness of algorithmic decisions Ghallab (2019). The disparate treatment of such models towards African Americans and women has been illustrated in the well-documented COMPAS Angwin et al. (2022) and Apple credit card Vigdor (2019) cases, respectively. While there has been extensive research on fair algorithms and fair machine learning, the proposed solutions have mostly disregarded important challenges that arise in real-world settings.
@@
Figure 1: Motivation: In Figure (a), a graph machine learning algorithm is applied to decide who receives credit (positive or negative) based on a possibly missing sensitive attribute, gender (and binary only for illustrative purposes). As shown in Figure (b), traditional missing data imputation does not account for outcomes (positive/negative), and thus, their imputed values can under-represent the bias of the complete dataset—demographic parity (DP) is $0.09$ in this example. This paper proposes BFtS, an imputation method for graph data that optimizes fairness in the worst-case imputation scenario using adversarial learning, as shown in Figure (c), where DP is $0.47$.
@@
Existing work in fair machine learning has mostly focused on tabular, image, and text data. However, in several applications, data can be naturally modeled as graphs (or networks), representing different objects, their relationships, and attributes. Graph Neural Networks (GNNs) have achieved state-of-the-art results in graph machine learning tasks, including node classification, link prediction, and graph classification Gilmer et al. (2017); Kipf and Welling (2017). For instance, in the case of link prediction in a professional network, such as LinkedIn, which is a recruiting and networking tool, link recommendations should not be biased against protected groups Fabbri et al. (2020); Current et al. (2022). However, guaranteeing fairness in graph data is a challenge due to correlations in the connections caused by homophily and influence.
@@
Another challenge for fair machine learning is the pervasiveness of missing data Little and Rubin (2019). This is particularly critical because fair algorithms often require knowledge of sensitive attributes that are more likely to be missing due to biases in the collection process or privacy concerns. As an example, the absence of data in census surveys and health surveys is correlated with sensitive attributes such as gender, age, race, etc. O’Hare and O’Hare (2019); Weber et al. (2021). Missing data is also a concern in networked data, especially when tracking the transmission of diseases Ghani et al. (1998). As a consequence, many fair machine learning algorithms require the use of missing imputation methods to mitigate their biases, and biases in imputed values might be inadvertently propagated to the final decisions Mansoor et al. (2022).
@@
This paper addresses the problem of mitigating bias in graph machine learning in the presence of missing sensitive attributes. We propose an adversarial imputation method that generates challenging instances for a broad class of fair algorithms based on GNNs. Figure 1 motivates our approach. A machine learning algorithm is applied to decide whether individuals should or should not receive credit. Figure 1(a) shows both the gender and outcome for each individual. The genders of some samples are unknown in Figure 1(a), and the demographic parity ($\Delta DP$) considering only the observed values is $0.25$ ($\Delta DP$ is a bias metric described in Section 4.2). We illustrate how different missing data imputations affect the fairness of credit decisions in Figures 1(b) and 1(c). A straightforward imputation is demonstrated in Figure 1(b), where the gender of the majority of neighbors is assigned to the missing attribute, resulting in a best-case scenario in terms of fairness with a $\Delta DP$ of $0.09$. However, in the worst-case scenario, shown in Figure 1(c), $\Delta DP$ is $0.47$. The dataset’s bias might be highly underrepresented in the imputation of Figure 1(b), and the final decisions of a fair algorithm will still be unfair in that case. We show the correlation of sensitive information and true class labels of different datasets (details in supp. material) in Figure 2. We use a Graph Convolutional Network (GCN) to predict the missing sensitive values. When $<50\%$ sensitive information is available, the predicted value underrepresented the true underlying bias in all datasets. Our method, BfTS, can better represent the true bias in the dataset. Our results provide strong evidence that off-the-shelf imputation methods often misguide fair graph machine learning algorithms due to underestimated bias.
Figure 2: Bias in training (i.e., the correlation between the sensitive attribute and node labels) for independent imputation and our approach (BFtS) with varying missing ratios. The rightmost column shows the true bias. Compared with BFtS, independent imputation often underestimates the true bias in the data.
@@
We propose Better Fair than Sorry (BFtS), which is a 3-player adversarial learning framework for missing sensitive value imputation. Figure 3 provides an overview of the proposed approach. The GNN $f_{C}$ interacts with two adversaries $f_{A}$ and $f_{I}$. $f_{A}$ attempts to predict sensitive values based on the representations learned by $f_{C}$ while $f_{I}$ imputes missing values that help $f_{A}$ maximize its accuracy. As a high accuracy of $f_{A}$ indicates high bias in the representations learned by $f_{C}$, $f_{C}$’s parameters are optimized to minimize the bias of its representations in the worst case in terms of fairness. Our experiments show that BFtS achieves a better accuracy $\times$ fairness trade-off than existing approaches. We summarize the contributions of this paper as follows:
@@
  • We propose Better Fair than Sorry (BFtS), a novel 3-player adversarial learning framework for the imputation of missing sensitive data that produces worst-case imputed values for fair GNNs and is effective even when sensitive attribute information is completely unavailable.
  • We prove that the maximal demographic parity (Eq. 8) of BFtS is minimal among all estimators of missing sensitive attributes and that BFtS is robust to convergence issues in adversarial training.
@@
  • We show empirically that BFtS achieves a better fairness $\times$ accuracy trade-off than the existing alternatives.
@@ 1.1 Related Work
Fairness in graph machine learning. We divide the work in this area into three main directions (pre-processing, extended objective functions, and adversarial learning). FairOT, FairDrop, FairSIN, EDITS, and Graphair are pre-processing techniques that remove the bias from the graph before training Laclau et al. (2021); Spinelli et al. (2021); Yang et al. (2024); Dong et al. (2022); Ling et al. (2023). Fairwalk, Crosswalk, Debayes, MONET, NIFTY, FairVGNN, PFR-AX, and FairGAE modify the objective function of GNNs to learn fair representations during training Rahman et al. (2019); Khajehnejad et al. (2022); Buyl and De Bie (2020); Palowitch and Perozzi (2019); Agarwal et al. (2021); Wang et al. (2022); Merchant and Castillo (2023); Fan et al. (2021). CFC, FLIP, DKGE, and Debias train adversarial networks along with GNN for fair prediction Bose and Hamilton (2019); Masrour et al. (2020); Arduini et al. (2020); Zhang et al. (2018). However, different from our work, all these methods require fully observed sensitive attributes to ensure fairness.
@@
Missing Data Imputation and fairness. Missing data can be imputed using unconditional mean, reconstruction, preferential attachment, autoencoders, etc. Donders et al. (2006); Pereira et al. (2020); Huisman (2009). The traditional procedure for handling missing data is to first impute the missing values, and then solve the task Rossi et al. (2021); Buck (1960). SAT is a matching-based GNN for graphs with missing attributes Chen et al. (2020), but it does not account for fairness. Training a classifier from imputed data can amplify the unfairness of a machine learning model, as discussed in Subramonian et al. (2022); Zhang and Long (2021); Feng et al. (2023). Some studies try to generate fair feature imputations. Subramonian et al. (2022) has modified mean aggregation feature imputation to minimize discrimination risk against protected groups. Feng et al. (2023) has proposed recursive partitioning and fair uniform resampling for fair classification with missing values. However, as the approaches discussed earlier, these methods require full knowledge of the protected attributes.
@@
Fairness with missing protected attributes. These studies are the ones most related to ours. One extreme assumption is the complete unavailability of sensitive information. Hashimoto et al. (2018) takes into account the worst-case distribution across a given group size, and Lahoti et al. (2020) adversarially reweighs training samples to optimize the average loss. Chai and Wang (2022) employs top $k$ average loss minimization to guarantee fairness. Zhao et al. (2022) minimizes the correlation between the predictions and other features associated with sensitive attributes. Yan et al. (2020) achieves fairness via class re-balancing and Chai et al. (2022) uses soft labels from an overfitted teacher model. In practice, some sensitive information is often available and can significantly improve fairness Chai and Wang (2022). FairGNN Dai and Wang (2021) assumes limited sensitive information and learns a fair GNN by predicting the missing values with an independent imputation model. RNF Du et al. (2021) generates proxy annotations for missing attributes using generalized cross-entropy Zhang and Sabuncu (2018). Similar to RNF, our approach (BFtS) handles completely or partially missing sensitive information. BFtS applies a 3-player scheme to minimize the maximum possible bias and outperforms FairGNN and RNF in terms of fairness $\times$ accuracy. Minimizing the maximum instead of the average risk has been shown to achieve better guarantees Shalev-Shwartz and Wexler (2016). A similar minmax approach has been applied to maximize the robustness and accuracy of uncertainty models Löfberg (2003); Lanckriet et al. (2002); Chen et al. (2017); Fauß et al. (2021). In Nguyen et al. (2017); Vandenhende et al. (2019), a 3-player adversarial network is proposed to improve the classification and stability of adversarial learning.
2 Problem Definition and Preliminaries
In this section, we formalize the problem addressed in this paper and introduce the Graph Neural Network model, which is our base model for node classification. For simplicity, we only consider binary-sensitive attributes. By modifying the loss function in Eq. 2 to categorical cross entropy and MSE, BFtS can be generalized for non-binary and continuous protected attributes, respectively.
2.1 Notation and Problem Definition
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X},\mathcal{S})$ be an undirected graph where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, $\mathcal{X}$ are node attributes, and $\mathcal{S}$ is the sensitive attribute. The matrix $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix of $\mathcal{G}$ where $A_{uv}=1$ if there is an edge between $u$ and $v$ and $A_{uv}=0$, otherwise. We focus on the setting where sensitive attributes might be missing and only $\mathcal{V}_{S}\subseteq\mathcal{V}$ nodes include the information of the sensitive attribute $s_{v}$ for a node $v$. The sensitive attribute forms two groups, which are often called sensitive ($s_{v}=1$) and non-sensitive ($s_{v}=0$).
We consider the fair node classification task, where the goal is to learn a classifier $f_{C}$ to predict node labels $y_{v}\in\{0,1\}$ based on a training set $\mathcal{V}_{L}\subseteq\mathcal{V}$. Without loss of generality, we assume that the class $y=1$ is the desired one (e.g., receive credit or bail). Given a classification loss $\mathcal{L}_{C}$ and a fairness loss $\mathcal{L}_{A}$, the goal is to minimize their combination:
$\min_{f_{C}}\mathcal{L}_{C}(f_{C})+\alpha\mathcal{L}_{A}(f_{C})$
Computing fairness loss $\mathcal{L}_{A}$ requires sensitive attributes $s_{v}$. The main challenge addressed in this paper is due to our assumption that sensitive attributes can be missing. In this scenario, one can apply an independent imputation model $s_{v}\approx f_{I}(X[v])$ to predict missing sensitive attributes before training the fair classifier. However, the fairness of the resulting classifier will be highly dependent on the accuracy of $f_{I}$, which is the key motivation for our work. In particular, we consider two scenarios to motivate our solution:
Sensitive attributes missing completely at random (MCR) at a high rate:
The missingness process can be modeled by a Bernoulli distribution $Bern(\beta)$ with $\beta$ close to $1$ and thus $f_{I}$ is expected to have a high error rate. Thus, the bias $\mathcal{L}_{A}$ can be underestimated by the fair classification model.
Adversarily missing sensitive attributes:
The missingness process is generated by an adversary that manipulates $f_{I}$ to artificially decrease the fairness loss $\mathcal{L}_{A}$. As an example, the adversary might select missing values to flip the predictions of $s_{v}$ for nodes that are both in the protected group and the undesired class ($y=0$). These scenarios are particularly relevant in graphs as they can exploit structural graph properties (see Sec. 4.1).
As an independent imputation cannot be trusted in these scenarios, we propose an interactive (3-player) worst-case imputation model for fair classification described in Sec. 3.
2.2 Graph Neural Network
Two of the models (or players) learned by our model are GNNs. Without loss of generality, we will assume that they are message-passing neural networks Gilmer et al. (2017), which we will simply refer to as GNNs. A GNN learns node representations $\mathbf{h}_{v}$ for each node $v$ that can be used to predict (categorical or continuous) values $y_{v}=f(\mathbf{h}_{v})$. These representations are learned via successive $\mathsf{AGGREGATE}$ and $\mathsf{COMBINE}$ operations defined as follows:
$\mathbf{h}_{v}^{k}\!=\!\mathsf{COMBINE}^{k}(\mathbf{h}_{v}^{k-1},\mathsf{AGGREGATE}^{k-1}({\mathbf{h}_{u}^{k-1},A_{u,v}\!=\!1})).$
Figure 3: Proposed 3-player framework for fair GNN training with missing data imputation (BFtS). $f_{C}$ generates node representations by minimizing the classification loss $\mathcal{L}_{C}$ (Eqn. 1) and maximizing sensitive attribute prediction loss $\mathcal{L}_{A}$ (Eqn. 3). $f_{A}$ predicts sensitive attributes using the representations from $f_{C}$ by minimizing $\mathcal{L}_{A}$. $f_{I}$ predicts missing values by minimizing the imputation loss $\mathcal{L}_{I}$ (Eqn. 2) and maximizing $\mathcal{L}_{A}$. $\hat{y}$, $\hat{si}$, $\hat{sa}$ are the predictions generated by $f_{C}$, $f_{I}$ and $f_{A}$, respectively.
@@ 3 Fairness-Aware Adversarial Missing Data Imputation for Graphs
We introduce a graph learning framework, which we call BFtS (Better Fair than Sorry), that allows fair GNN training in a setting with missing sensitive data. BFtS applies a 3-player adversarial learning scheme, where the fair GNN is trained jointly with two adversaries—one to predict sensitive attributes based on GNN embeddings and another to impute missing values that minimize fairness. As a result, fairness assessments by the GNN are based on the worst-case imputed values generated by the adversarial imputation.